OpenStudy (anonymous):

Find the inverse function of F(x)=[(2x+1)/(x-3)] I'm having trouble with the algebra steps in between

OpenStudy (cwrw238):

let y = 2x+1 / x - 3 xy - 3y = 2x + 1 xy - 2x = 3y + 1 x = 3y - 1 / y - 2 inverse = (3x -1) / (x - 2)